### v3

# Superconductivity in overdoped cuprates can be understood from a BCS perspective!

Abstract
We summarize key experimental studies of the low energy properties of overdoped cuprate high temperature superconductors and conclude that a theoretical understanding of the “essential physics” is achievable in terms of a conventional Fermi-liquid treatment of the normal state, and a BCS mean-field treatment of the ($d$-wave) superconducting state. For this perspective to be consistent, it is necessary to posit that there is a crossover from a strongly correlated underdoped regime (where a different theoretical perspective is necessary) to the more weakly correlated overdoped regime. It is also necessary to argue that the various observed features of the overdoped materials that are inconsistent with this perspective can be attributed to the expected effects of the intrinsic disorder associated with most of the materials being solid state solutions (alloys). As a test of this idea, we make a series of falsifiable predictions concerning the expected behavior of an “ideal” (disorder free) overdoped cuprate.
I Introduction
Background: The search for a theoretical understanding of the experimentally documented properties of cuprate high temperature superconductors has played a central role in condensed matter physics for almost four decades, in large part because these materials have been more intensively and incisively studied than any other correlated electronic material. Influential and broadly relevant ideas that germinated from such studies include the physics of quantum antiferromangets (including possible spin-liquids) and doped antiferromagnets, novel forms of charge, spin, current, and superconducting density wave orders, electronic analogues of various liquid crystalline phases, ways in which usual notions of metallic physics can break-down (non-Fermi liquid behaviors), bold new approaches to obtaining controlled (numerical) solutions of model problems - such as the Hubbard model - and a host of more phenomenological ideas focused on various experimental observations. Indeed, the observation of multiple novel behaviors in these materials justifies this exuberant exploration of various interesting observed phenomena.
One notable feature of this field is the enormous effort that has been devoted to obtaining exquisite control over the materials involved. Results are now highly reproducible. However, one key feature we will stress is that this does not mean that the effects of “disorder” can be neglected. Most of the materials involved are alloys, meaning that even in “perfect” materials, there is inherent disorder associated with the random configuration of dopant atoms. While we are certainly interested in the measured properties of the actual materials, it is also useful to ask what the features of a putative “ideal” (i.e. disorder free) cuprate would be.
A central set of questions in the field concern the nature of the superconducting state itself: Why is $T_{c}$ so high? What limits $T_{c}$? Why do the phase diagrams of many families of cuprates exhibit a relatively similar superconducting “dome?” What do we learn from the cuprates about unconventional superconductivity in other material systems?
Foreground: In this article (which is a perspective, and not a review), we analyze the evidence that the superconducting dome in an “ideal” disorder free cuprate would arise from a crossover between a strongly correlated (and hence very complex) underdoped regime, where $T_{c}$ is intrinsically determined by phase ordering, to an overdoped regime, where it is possible to adopt a weak coupling “BCS” perspective based on a Fermi liquid normal state and a mean-field description of the superconducting state (but not, necessarily, invoking phonon-mediated pairing). As part of this analysis, we present reasons to attribute to the effects of disorder those observed features of actual overdoped cuprates that are inconsistent with a weak-coupling description. Our perspective is summarized in Figure 1, for which the essential assumptions are:
  • $p<p_{opt}$: There is an underdoped (UD) regime that may be best thought of as a lightly doped antiferromagnetic insulator. Here, the pairing scale $|\Delta_{0}|$ is large compared to the superconducting $T_{c}$. Conversely, the energy scale that characterizes the stiffness to phase fluctuations,[1] $T_{\theta}$, is small compared to $|\Delta_{0}|/2$. Consequently, $T_{c}$ is the temperature at which phase fluctuations, rather than thermal pair-breaking, destroy long-range order. This hierarchy of scales, $|\Delta_{0}|/2\gg T_{c}\approx T_{\theta}$, is inverted from the usual weak-coupling case, and so likely reflects strong-coupling physics. Indeed, in this regime of doping there is direct evidence for various types of “intertwined” density-wave-orders (which are generically strong-coupling effects) and a host of fluctuation phenomena which have been linked under the name “pseudo-gap phenomena.” There is no generally accepted theory that describes the rich physics of the underdoped regime.
    Figure 1: Conjectured schematic phase diagram of an ideal disorder-free hole-doped cuprate. $T_{\Theta}$ is determined by $\kappa$ the zero-temperature superfluid stiffness ( $\kappa_{0}\sim\hbar^{2}n_{s}/m^{\star}$), and $c$, the mean distance between Cu-O planes (so $\kappa\ c$ is the 2D superfluid stiffness per plane). $T_{\Delta}$ is determined by the magnitude of the $T=0$ superconducting (pairing) gap, $\Delta_{0}$. $T_{\rm c}$, shown in black, is set by the lower of these two scales. The red shaded region indicates a crossover from an effective Fermi-liquid-like overdoped regime to a highly correlated underdoped regime, where the normal state is not well described in terms of weakly interacting quasi-particles, and where various forms of intertwined order further complicate matters. At low doping, the blue region contains various phases with (at low $T$) long range antiferromagnetic order - either commensurate or incommensurate - and various forms of hole crystals, ranging from hole-Wigner crystals at the lowest doping to stripe-ordered insulators at higher doping. Aspects of this conjectured phase diagram are inconsistent with experiments, as can be seen, for example, in Figs. 2 and 3; we attribute these discrepancies to the effects of disorder.
  • $p>p_{opt}$: On the overdoped (OD) side of the superconducting dome, Figure 1 shows $T_{\theta}\gg T_{c}\approx|\Delta_{0}|/2$, the familiar hierarchy present in BCS superconductors. Confirming this behavior would constitute strong evidence of the applicability of a weak-coupling perspective. Correspondingly, the normal state, with superconductivity suppressed either with increasing temperature or magnetic field, should be well captured by the usual Fermi liquid theory of metals.
  • $p\sim p_{opt}$: For a range of dopings in the vicinity of optimal doping, $p=p_{opt}$, there is a regime where both $|\Delta_{0}|/2$ and $T_{\theta}$ are comparable to $T_{c}$. This is a crossover regime, where presumably no single perspective can easily capture all of the physics.
The main point: Our primary purpose is to critically asses the validity of a weak coupling perspective as it relates to the overdoped cuprates. (While most theoretical studies of the cuprates have emphasized more exotic, strong coupling aspects of the problem, there is a significant subset that overlap either in some details or in perspective with the present approach [2, 3, 4, 5, 6, 7, 8, 9, 10, 11].) Specifically, we identify the experimental evidence that supports this perspective and, perhaps more importantly, observations that clearly contradict it. For example, several families of cuprates exhibit the “boomerang effect,” where $T_{\theta}$ decreases with increasing $p$ such that it remains comparable to $T_{c}$ throughout the overdoped regime (thus appearing to “turn around” when plotted parametrically as a function of $p$ on a $T_{\theta}$ vs $T_{\rm c}$ plot)[12, 13, 14, 15]. Moreover, when superconductivity is suppressed with a high magnetic field, “strange metal” behavior (as opposed to the expected Fermi-liquid behavior) is seen throughout the overdoped range of $p$ [16, 17]. We will discuss reasons to think that these features—those that challenge our perspective when taken at face value—are the consequence of intrinsic alloy-disorder.
Along these lines, we present evidence that the apparent failures of the weak-coupling perspective are less pronounced in materials in which such intrinsic disorder is arguably weaker. The sketched phase diagram moreover stands as a falsifiable prediction - if experiments are carried out in materials with less intrinsic disorder (e.g. stoichiometric crystals) and they fail to produce results that are increasingly consistent with the sketched phase diagram, then the underlying theory cannot be correct.
The conceptual underpinnings of the perspective presented here are distinct from those underlying the “two-fluid” phenomenology (2FP) discussed in Ayres et al. [18]. Whereas we identify as “extrinsic” (i.e. vanishing in the zero disorder limit) the various properties of overdoped materials that are inconsistent with a FL/BCS description, the 2FP interprets them as consequences of unspecified strong correlation effects that lead intrinsically to the existence of coexisting “coherent” and “incoherent” fluids of electronic excitations.
The present perspective is also fundamentally different from the widely discussed perspective in which optimal doping is associated with a quantum critical point whose influence dominates large portions of the phase diagram.[19, 20, 21] While our perspective does not preclude the existence of one or more quantum critical points with $p_{c}\lesssim p_{opt}$, one outcome of our discussion will be to adduce evidence—under the assumption that the weak-coupling analysis is reasonable for $p>p_{opt}$—that the pairing interaction does not support the idea that a putative quantum critical point at $p^{*}\sim p_{opt}$ plays a significant role in the pairing mechanism. (see the supplementary information for further discussion of the role of quantum criticality).
There is a lot riding on this. Admittedly, without addressing the crossover to the strongly correlated physics on the UD side of the phase diagram, we cannot claim to fully understand superconductivity at optimal doping, nor why $T_{c}$ drops with underdoping. However, if the present perspective is verified, it gives us an understanding of high temperature superconductivity from the OD side. As the cuprates are the best characterized of all unconventional SCs, even this partial understanding has broad implications for the field.
From the weak-coupling perspective, there are two key features of the cuprates: i) a band-structure with a large density of states near the $(0,\pi)$ and $(\pi,0)$ points in the BZ, and ii) a $\vec{k}$ dependent repulsive interaction peaked near $\vec{Q}=(\pi,\pi)$, i.e. the vector that connects the high density of states regions. (The presence of this interaction is, in turn, reflected in the existence of pronounced short-range magnetic correlations with roughly the same wavevector.) This invites us to search for new high temperature superconductors by looking for band-structure motifs and magnetic correlations that remind us of the cuprates, rather than trying to reproduce the local solid state chemistry.
II What is a BCS superconductor?
Using the words “BCS” and “cuprate” in the same sentence is obviously provocative. As elaborated below, we apply the term ‘BCS superconductor’ to a situation in which the superconducting state emerges as a weak-coupling instability from a Fermi liquid normal state, independent of whether phonons play a role in the mechanism or not.
In what sense are “conventional” superconductors described by BCS theory?
The properties of crystalline (“clean”) conventional superconductors can be derived from the electron-phonon problem using the Migdal-Eliashberg (ME) approximation. Unlike the original BCS theory, which abstracts away the pairing mechanism with a generic attractive potential, ME explicitly keeps both the electrons and the “bosonic glue,” i.e. phonons.
At temperatures $T>\hbar\omega_{D}$ (where $\omega_{D}$ a characteristic phonon frequency - e.g. the Debye frequency), the quasi-particle lifetime (in both theory and experiment) is $\hbar/\tau=2\pi\lambda_{0}T$, where $\lambda_{0}$ is the dimensionless (unrenormalized) electron-phonon coupling constant. It is not immediately obvious that Fermi liquid theory applies at these temperatures, where quasiparticles (in the Landau sense) are at best marginally well defined [22]. However, if one is only interested in treating the problem at temperatures, $T$, and energies, $\hbar\omega$, small compared to $\hbar\omega_{D}$, then it is possible to introduce (as was done first by BCS) a simpler low-energy description of the system in terms of well defined Landau quasi-particles with a weak residual interaction proportional to the dimensionless parameter, $\lambda-\mu^{\star}$. Here, $\lambda\sim\lambda_{0}/(1+\lambda_{0})$, and $\mu^{*}$ are renormalized versions of the dimensionless electron-phonon and electron-electron interactions. Formally, this “effective field theory” can be derived[23, 24, 25] from the original problem by integrating out degrees of freedom with energies larger than $\sim\hbar\omega_{D}$, so that $\hbar\omega_{D}$ appears only as the high energy cuttoff of the theory.  [^1] There are a few subtleties concerning the BCS perscription for integrating out the phonons that are worth mentioning: i) As there are always acoustic phonons, it is not literally the case that all the phonons get integrated out in this process. However, these modes do not couple strongly to the electrons - both because of limited phase space and the fact that, being Goldstone modes, they only have derivative couplings to electrons. They can thus be neglected. ii) Even at low $T$, there can be dynamical effects - such as relaxation of non-equilibrium populations of quasi-particles - that cannot be well understood in terms of the effective field theory, but require a theory in which phonons are included explicitly.
It is interesting to note that an essential piece of the evidence that the “mechanism” of pairing is a phonon-induced attraction between electrons comes from analyzing the energy dependence of the gap function, $\Delta(\vec{k},\omega)$, at energy scales $\hbar\omega\sim\hbar\omega_{D}$, i.e. at energies where Landau quasiparticles are not sharply defined. In the context of the upcoming discussion of the overdoped cuprates, a take-away message is that the lack of a conventional quasiparticle description at energies much greater than $T_{\rm c}$ does not preclude the applicability of a weak-coupling BCS perspective at low energies. To gain additional insight, let us look at a conventional superconductor from a low $T$, low energy empirical perspective and imagine trying to infer something about the pairing mechanism. To begin with, we would note that if we suppress superconductivity with a magnetic field, we would see well-defined quantum oscillations (QOs) - telling us that the normal state is a Fermi liquid metal. From the facts that the gap vanishes at $T_{c}$, that $T_{c}$ is (up to corrections that are never more than 20-30% effects) is equal to $|\Delta_{0}|/2$, and that thermodynamic quantities, such as the specific heat, show mean-field like behavior, we would be confident that a mean-field description is warranted. Finally, once we succeeded in measuring the gap function, $\Delta(\vec{k},\omega)$, we would see that it is only weakly $\vec{k}$ dependent: from this we would infer that pairing is induced by an effective attractive interaction that is local (short-ranged) in space. That is to say, we would identify phenomena that are consistent with the underlying assumptions of BCS theory.
More detailed information about the pairing mechanism is somewhat hidden in this BCS approach. In principle, the attractive interaction could come from something other than phonons. Indeed, this is the sense in which we are suggesting that overdoped cuprates can be described by BCS theory: it is clear that the pairing interaction in the cuprates is not, primarily, due to phonons, and yet we will argue that a BCS perspective is warranted.  [^2] It is also not entirely true that there are no signatures of the mechanism; in conventional SCs, the importance of phonons can (often) be derived from the isotope dependence of the cutoff in the theory, an observation that played an essential role in the acceptance of BCS theory.
When are fluctuations negligible?
BCS theory is a mean-field theory and therefore is only valid so long as fluctuation effects are negligible. In principle, close enough to $T_{c}$, universal critical fluctuations must arise, meaning that BCS theory is never applicable. However, in conventional superconductors, significant fluctuation effects are confined to extremely narrow regimes about $T_{c}$ (of order $10^{-8}$$T_{\rm c}$ [26]).
While there are multiple possible sorts of fluctuation effects that can give corrections to BCS theory, particularly important in the cuprate context are “phase fluctuations.” In the superconducting state below $T_{c}$, the low energy physics is universally captured by the non-linear sigma model,
$F=\frac{\kappa(T)}{2}\ \int d\vec{r}\ \left|\vec{\nabla}\theta-\frac{2e}{\hbar c}\vec{A}\right|^{2}$ (1)
where $\kappa$ is known as the superfluid stiffness, $\theta(\vec{r})$ is the local phase of the superconducting order parameter, and $\vec{A}$ is the electromagnetic vector potential. Importantly, $\kappa$ is a measurable quantity: in 3d, $\kappa(T)=[(\hbar c)^{2}/16\pi e^{2}]\ \lambda^{-2}(T)$ where $\lambda$ is the London penetration depth. We define an energy scale associated with phase fluctuations as
$k_{\rm B}T_{\theta}(T)\equiv\kappa(T)L^{d-2}$ (2)
where $L$ is an appropriate microscopic length scale that provides a UV cutoff to the theory. (We will adopt the notation $T_{\theta}\equiv T_{\theta}(T=0)$.) Of particular interest here is that in $d=2$, $\kappa$ has units of energy. For a layered system of weakly Josephson coupled superconducting planes, the appropriate length scale is $L=c$, i.e. the spacing between planes. (More generally, $L$ is the larger of $c$ and the superconducting coherence length, $\xi_{c}$, in the interplane direction.)  [^3] One consequence of this is that close enough to $T_{c}$ the physics must universally crossover to 3D criticality, in which case $L\sim\xi_{c}$.
The meaning of this energy scale is that thermal phase fluctuations are dilute (and hence in certain senses negligible) so long as the temperature is much less than $T_{\theta}(T)$, while they have an order one effect on the physics when the temperature is of order $T_{\theta}(T)$ or greater. Close enough to $T_{c}$, phase fluctuations are always important: In 2D, as $T\to T_{c}$ from below, $T_{\theta}(T)\to(2/\pi)T_{c}$ (i.e. there is a universal drop of the phase stiffness at $T_{c}$). In 3D (to the extent that gauge-field fluctuations can be ignored), $T_{\theta}$ goes to zero as $T\to T_{\rm c}^{-}$, with universal 3D XY critical exponents. However, in both cases, if at low temperatures $T_{\theta}\gg T_{c}$, it is only in a range of temperatures of order $\Delta T\sim T_{c}(T_{c}/T_{\theta})$ that phase fluctuations are expected to be important. Thus, even in the case of a film of a conventional clean superconductor of microscopic thickness $W$, since $T_{\theta}\sim E_{F}(k_{F}W)\gg T_{c}$, universal 2D XY behavior is essentially unobservable.
III What would a BCS approach to the cuprates entail?
BCS approach to the overdoped cuprates
Because the cuprates are strongly correlated, the theoretical route from the microscopic physics of copper and oxygen orbitals to an appropriate low energy effective model is difficult and will not be addressed here. Instead, we assume that there exists an intermediate energy scale that is large compared to $T_{c}$—possibly of order the exchange energy in the parent antiferromagnetic insulator, $J\approx 1500$ K [27]—below which the electronic properties can be understood on the basis of an effective model of well defined quasi-particles with relatively weak residual interactions. The quasi-particle properties may be strongly renormalized, and the residual interactions can encode the effects of important high energy physics, especially the existence of local moments on the copper sites. Importantly, however, because the effective theory is obtained by integrating out high energy degrees of freedom, the residual interactions can be taken to be short-ranged, and both the interactions and the quasi-particle renormalizations must be smooth functions of microscopic parameters, such as the doping concentration, $p$. The existence of quantum oscillations in some overdoped cuprates, to be discussed in more detail later, constitutes strong evidence that this assumption is reasonable, at least at low temperatures.
The BCS self-consistency condition is derivable as an implicit expression for the gap function, $\Delta(\vec{k})$ (which has no explicit $\omega$ dependence because we are dealing with an effective Hamiltonian), as a function of the effective interaction, $\tilde{V}(\vec{r})$. This condition is usually expressed as a convolution over momentum, $\vec{q}$, of $V(\vec{q})$, the Fourier transform of $\tilde{V}(\vec{r})$, and an integrand that depends on the band dispersion, $\epsilon(\vec{k})$ and the quasi-particle energy,
$E(\vec{k})\equiv\sqrt{|\epsilon(\vec{k})-\mu|^{2}+|\Delta(\vec{k})|^{2}},$ (3)
where $\mu$ is the chemical potential. However, it is also useful to look at this relation in real-space:
$\tilde{\Delta}(\vec{r})=-\tilde{V}(\vec{r})\ \tilde{\phi}(\vec{r}),$ (4)
where $\tilde{\phi}(\vec{r})=(1/2)\langle\left[c_{\vec{r},\uparrow}^{\dagger}c_{\vec{0},\downarrow}^{\dagger}+c_{\vec{0},\uparrow}^{\dagger}c_{\vec{r},\downarrow}^{\dagger}\right]\rangle$ is the pair wave-function,
$\tilde{\phi}(\vec{r})=\int\frac{d\vec{k}}{(2\pi)^{d}}\ e^{i\vec{k}\cdot\vec{r}}\ \frac{\Delta(\vec{k})}{2E(\vec{k})}\tanh\left[\frac{E(\vec{k})}{2T}\right]\ F\left[\epsilon(\vec{k})\right],$ (5)
where $F(\epsilon)$ implements a high energy cutoff of the theory, if needed.
One feature of BCS theory that is often overlooked is that the form of the gap function is almost entirely determined by the form of the effective interaction. This follows from Eq. 4 and the fact that $\tilde{\phi}$ is much longer ranged (more slowly varying as a function of $|\vec{r}|$) than $\tilde{V}$. This explains why the gap function $\Delta(\vec{k})$ is typically approximately momentum independent in conventional superconductors. It also accounts for the fact that, although the magnitude of $\Delta(\vec{k})$ changes with temperature and vanishes at $T_{c}$, its $\vec{k}$ dependence remains roughly constant—even in unconventional superconductors. For the case of a $d$-wave SC, relevant to the cuprates,
$\Delta(\vec{k})=\Delta_{0}\ d(\vec{k})\ f(\vec{k})$ (6)
where $\Delta_{0}$ characterizes the magnitude of the gap function, $d(\vec{k})\equiv(1/2)[\cos(k_{x})-\cos(k_{y})]$ is mandated by the $d$-wave symmetry, and $f(\vec{k})$ (which determines the range of $\tilde{\Delta}$) transforms trivially under all point group transformations and is normalized so that $f(0,\pi)=f(\pi,0)=1$.
In the case of the shortest-range interactions allowed by symmetry, $f(\vec{k})=1$, corresponding to a gap function that is non-zero only on nearest-neighbor bonds. To be explicit, we will treat the case of a density-density interaction, $\tilde{U}(r)$, and a spin-spin interaction (exchange coupling), $\tilde{J}(r)$, which vanish beyond nearest-neighbor distance: $\tilde{U}(r)=\tilde{J}(r)=0$ for $r>1$. In this case, whenever a non-zero $\Delta_{0}$ exists, the BCS self-consistency relation can be written as
$1=-\tilde{V}\int\frac{d\vec{k}}{(2\pi)^{d}}\ \frac{|d(\vec{k})|^{2}}{2E(\vec{k})}\tanh\left[\frac{E(\vec{k})}{2T}\right]\ F\left[\epsilon(\vec{k})\right]$ (7)
with $\tilde{V}=-\left[3\tilde{J}(1)-4\tilde{U}(1)\right]$. (Because we have assumed $d$-wave symmetry, the onsite interaction $\tilde{U}(0)$ does not play any role in pairing.) This equation applies only for the case $\tilde{V}<0$, and below $T_{c}$. Note, that while we consider it likely that the dominant contribution to $\tilde{V}$ is associated with incipient (short-range) antiferromagnetic interactions ($\tilde{J}(1)>0$), from the present perspective this cannot be distinguished from the effects of a short-range density-density attraction ($\tilde{U}(1)<0$) that could arise from interactions with appropriate high-energy optical phonon modes. Equation 7 is typically viewed from a bottom-up perspective as a route to compute $\Delta_{0}(T)$ from a given (assumed known) effective interaction; below, we will invert this logic, and use measured values of $\Delta$ to infer the nature of the effective interaction, $\tilde{V}$.
Within BCS theory, $T_{c}$ vanishes only when $\tilde{V}$ becomes positive. Since the dependence of the effective parameters on microscopic parameters is analytic by assumption, in the present context this means that if there is (in the absence of disorder) a critical value of the doping, $p_{max}$, beyond which $T_{c}$ vanishes, then
$\tilde{V}\approx-\tilde{V}_{0}\ (p_{max}-p),$ (8)
where $\tilde{V}_{0}>0$. Consequently, $T_{c}\sim e^{-\alpha(p_{max}-p)^{-1}}$ where $\alpha\propto|1/\tilde{V}_{0}|$ is a dimensionless constant. In other words, if the pairing interaction changes sign at a critical doping, $T_{\rm c}$ will decrease exponentially on approach to that doping, rather than cutting off sharply as it does at the edge of a “dome”. This is the basis of the overdoped “tail” shown in Figure 1.
Effects of disorder on a BCS d-wave superconductor
One might naively have thought that superconductivity in the weak coupling (BCS) limit would require extremely clean systems so that the mean-free-path, $\ell$, is large compared to clean limit coherence length, $\xi_{0}$. It is a special property of conventional s-wave SC, known as “Anderson’s theorem,” that both $T_{c}$ and $\Delta_{0}$ are largely independent of $\xi_{0}/\ell$. Generically, disorder only significantly affects $T_{c}$ when it is strong, i.e. when $k_{F}\ell\sim 1$.
However, it follows straightforwardly from the effective field-theoretic approach of Abrikosov and Gorkov (AG) that for $d$-wave and other superconductors with a sign-changing gap, the pair-breaking effect of disorder results in the elimination of the superconducting state when $\xi_{0}/\ell$ exceeds a critical value of order 1. Indeed, there are some features of the resulting theory that resemble observed properties of overdoped cuprates.[28, 29, 30] For one thing, AG theory gives an explicit expression for the way $T_{c}$ is suppressed with increasing $\xi_{0}/\ell$ in which $T_{c}$ vanishes linearly as $\xi_{0}/\ell$ approaches the critical value. Moreover, the zero temperature superfluid stiffness, $\kappa$, vanishes in proportion to $T_{c}$ as $T_{c}\to 0$. However, this does not indicate that phase fluctuations become important (and BCS theory therefore inaccurate). Specifically, in applying in Eq. 2 to 3D materials, one expects the length scale, $L$, to be the $c$-axis coherence length, $\xi_{c}$. Because $\xi_{c}$ diverges as $T_{c}\to 0$, it follows that $T_{\theta}/T_{c}$ also diverges. Even it we fix $L=c$,
$T_{\theta}/T_{c}\sim(k_{F}\ell)\sim(k_{F}\xi_{0})\gg 1\ {\rm as}\ T_{c}\to 0.$ (9)
As far as we know, accessing a regime in which $T_{\theta}/T_{c}\lesssim 1$ requires something beyond BCS theory.
AG theory treats disorder in an effective medium approximation. This is expected to be extremely accurate in the usual situation in which $\xi_{0}\gg L_{dis}$, where $L_{dis}$ is the correlation length of the disorder, a hierachy that applies to many unconventional superconductors including the iconic example of Sr2RuO4.[31]
Even in the context of BCS mean-field theory, the cuprates are special both in that they are d-wave and that they have a relatively small coherence length, $\xi_{0}\sim$ 2 or 3 lattice constants. In particular, $\xi_{0}$ is comparable to $L_{dis}$, which is likely of order $p^{-1/2}$ lattice constants;  [^4] As has been stressed by Tranquada,[32] the RMS variation of the dopant concentration in a region of size $L$ is $\delta p\sim\sqrt{p}/L$. Moreover, because the dopant atoms are generally located away from the Cu-O plane, and affect the electronic structure in these planes both through Coulomb interactions and induced strain, even if the dopant positions are entirely uncorrelated the disorder potential is expected to have spatial correlations of substantial extent. consequently, mesoscopic fluctuations in the local value of $\ell$ inevitably lead to large variations in the local value of $T_{c}$. Therefore, near the edge of the superconducting dome the effects of disorder are expected to be amplified, leading to an effectively granular state with locally superconducting puddles emersed in a metallic background. Various numerical and analytic studies have confirmed this expectation.[33, 6, 5, 34] Importantly, the inevitable consequence is a more rapid suppression of $T_{\theta}$ than of the spatially averaged $T_{c}$. Thus, near the edge of the superconducting dome, phase fluctuations inevitably become important and lead to a breakdown of the mean-field description.
III.1 Additional considerations
In the Supplemental Materials, we expand upon certain important matters of perspective. In Sec. SM-I, we review the difference between a crossover and a phase transition in the context of the assumed crossover indicated by the rose region in Fig. 1. This is essential to the present analysis in that it justifies treating the mechanism of superconductivity in the overdoped regime without needing to account for the behavior of the underdoped regime, despite the fact that they are smoothly (adiabatically) connected to one another. In Sec. SM-II, we discuss aspects of the relation between the present perspective and perspectives that treat the existence of a putative quantum critical point at $p=p^{\star}$ under the superconducting dome as the central organizing feature of the cuprate phase diagram. In Sec. SM-III, we expand upon the tangible differences between homogeneous “dirty $d$-wave” and inhomogeneous, self-organized granularity as the cause of the suppressed superfluid density on the overdoped side of the phase diagram.
IV Experiments on overdoped cuprates
Here, we review a body of experimental evidence that contains useful information concerning the applicability of the BCS perspective in overdoped cuprates. We summarize sets of experimental observations that most clearly support the proposed perspective and then those that are inconsistent with or apparently inconsistent with it. In each case, we present arguments with features that are inconsistent with the BCS perspective can be plausibly understood as fluctuation effects arising from the sort of inhomogeneous electronic structure expected for a disordered $d$-wave superconducting with a short coherence length - i.e. that in the absence of disorder the features that are inconsistent with BCS theory would be absent.
The body of experiments discussed here is necessarily much smaller than the total corpus. 1) We focus on experiments relevant to overdoped cuprates, and only touch on underdoped experiments where the comparison is helpful.  [^5] To further focus the discussion, we consider only hole-doped cuprates. As mentioned below, the significant antiferromagnetic correlations in the electron doped cuprates that persist to optimal doping and beyond, may complicate the present analysis applied to these materials. For a review, see [35]. 2) We focus on what we consider to be the “strongest” evidence: for example, we use quantum oscillations to argue for a Fermi liquid ground state rather than power laws in various transport coefficients. 3) We restrict our attention to “low energy” properties, that is to say those that reflect excitations with energies less than an intermediate scale which we take somewhat arbitrarily to be $\sim 0.1$ eV, or in other words a scale of order the exchange energy $J$ in the undoped parent materials. 4) We focus on a subset of materials: YBa2Cu3O6+δ (YBCO), the electronically cleanest cuprate; La2-xSrxCuO4 (LSCO), the cuprate with the most comprehensive data sets, including in the low-temperature normal state accessed by magnetic fields; Tl2Ba2CuO6+δ (Tl2201), the cleanest cuprate that can access the overdoped regime; and Bi2Ba2CaCu2O8+δ (Bi2212), which has the best angle-resolved photoemission data.
One issue that arises when comparing results reported in different studies on multiple families of cuprates is that the doped hole concentration, $p$, is typically inferred rather than measured. Indeed, often it is estimated on the basis of an assumed universal form of the $T_{c}(p)$ dome. Especially since we will conclude that the termination of the dome is highly sensitive to the character and nature of the disorder, we have not adopted this approach. Instead, we adopt (wherever possible) the value obtained from the area enclosed by the Fermi surface obtained from low energy angle-resolved photoemission spectroscopy (ARPES) measurements or from quantum oscillations.
The Fermi liquid ground state
Overdoped cuprates have quantum oscillations, angle-dependent magnetoresistance, and ARPES that imply a Fermi surface similar to what is found in local-density approximation (LDA) calculations.
Experiments in support of a Fermi liquid ground state.
Quantum oscillations (QOs)—arising due to Landau quantization of quasparticle orbits in a magnetic field—are the gold standard for demonstrating a Fermi liquid ground state. At high enough magnetic fields to quench superconductivity, QOs are observed in Tl2201 at dopings higher than $p=0.27$, corresponding to $T_{\rm c}^{\prime}s\leq 26$ K [36, 37]. QOs are not observed for $p<0.27$, possibly due to the presence of 2D short-range correlated CDW order,[38, 39] as discussed below.
Beyond the fact that there are QOs, which demonstrates the presence of a Fermi surface, we also learn:
  • There is a single, hole-like, cylindrical Fermi surface. The measured areas correspond to hole densities of $p=0.27$ for $T_{\rm c}=26$ K, $p=0.3$ for $T_{\rm c}=10$ K [36], respectively, and are consistent with those inferred from angle-dependent magnetoresistance [40], and ARPES [41].
  • The temperature dependence of the QOs fits the Fermi-liquid “Lifshitz-Kosevich” (LK) form very well from $0.3$ K to $3$ K, yielding a cyclotron effective mass of approximately $m^{\star}=5$ $m_{e}$, where $m_{e}$ is the bare electron mass.
  • The quasiparticle scattering rate is $\hbar/\tau=2.4$ meV. (Note that this is the not the transport scattering rate—QOs are sensitive to forward scattering processes that do not affect the resistivity).
  • The angle dependence of the quantum oscillations is consistent with a $g$-factor of 22 [36], similar to underdoped YBCO [42].
While the Fermi surface geometry agrees very well with LDA [43], a bandwidth renormalization of a factor of 3 is required to match the cyclotron mass with that calculated from LDA. This is similar to the mass renormalizations found in Sr2RuO4 [44]. The same bandwidth renormalization is found over hundreds of meV by ARPES [45]. As pointed out by Rourke et al. [36], this demonstrates that the quasiparticle mass renormalization arises predominantly through eV-scale electron-electron interactions, and not from interactions with a low-energy boson. Moreover, interpreting the temperature dependence of the QO amplitude as an effective mass assumes that corrections from an energy-dependent self-energy are small [46]. The absence of any deviation from the standard LK form in Tl2201 implies that there are no strongly energy-dependent renormalizations (as there would be in the marginal Fermi liquid [47]), even with a $T_{\rm c}$ as high as 26 K.
Experiments challenging a Fermi liquid ground state.
The strongest arguments against a Fermi liquid ground state stem from observations of various transport anomalies at low $T$ when superconductivity is suppressed by a high magnetic field. Of these, the most dramatic is the $T$-linear resistivity extending to low temperatures in a range of doping from $p_{opt}$ to the OD edge of the superconducting dome, $p_{max}$. [16]. (Note that here we are not discussing the $T$-linear resistivity extending to very high temperatures observed near optimal doping that has long been argued to indicate the existence of a non-Fermi-liquid ground state at or near a special doping, $p^{*}$.)
Specifically, Cooper et al. [16] report $T$-linear resistivity out to the very edge of the superconducting dome in LSCO—when superconductivity disappears, so does the $T$-linear resistivity. These measurements are consistent with earlier observations in Tl2201 [17]. It has also been suggested that the approximate correspondence between the ranges of $p$ in which they are observed implies that the scattering responsible for $T$-linear resistivity is also responsible for superconducting pairing. From a somewhat different perspective, the observation of $B$-linear magnetoresistance near optimal doping in LSCO [48], and later to higher dopings in Tl2201 and Bi2201 [18], and has been interpreted as evidence for “scale invariance”, as might be expected near a quantum critical point.
Reconciling a Fermi liquid ground state.
There have been several theoretical proposals attempting to reconcile the unambiguous evidence from QO’s for the existence of a Fermi liquid with the observed transport anomalies. As discussed above, even a high-quality, “uniformly doped” sample will appear heterogeneous on short enough length scales due to the statistical distribution of dopant locations. This can lead to an electronically inhomogeneous state that consists of dilute superconducting “puddles” that might survive even to the high fields at which superconducting long-range coherence is destroyed; a recent paper[49] has demonstrated that scattering from the quantum phase fluctuations on such puddles can give rise to a $T$-linear resistivity. Experimental evidence for such a scenario comes from overdoped Tl2201, where increasing the magnetic field from 10 to 18 tesla reduces the $T$-linear slope by at least a factor of 2 [50]. Alternatively, the possibility that anomalous scattering might be associated with local, nearly-quantum-critical patches of some other order - possibly associated with a disorder broadened antiferromagnetic quantum critical point - has been proposed in a series of related papers.[51, 52, 53, 54, 55]
For the present purposes, what is essential is that these proposals imply that an extrinsic (disorder related) form of anomalous scattering is responsible for the non-Fermi liquid transport. Our specific prediction is that the $T$-linear term will be absent in less disordered overdoped cuprates, such as stoichiometric YBCO doped through application of hydrostatic pressure [56].
Negligibility of superconducting fluctuations
When the phase ordering scale, $T_{\theta}$, is comparable to $T_{c}$, phase fluctuations necessarily are an order-one effect; this implies a breakdown of BCS mean-field theory. This situation applies in all underdoped cuprates. The issue is more complex in the overdoped cuprates.
Experimental evidence that fluctuations are insignificant
While in LSCO and Tl1201, the superfluid density (or more precisely $T_{\theta}$) decreases with increasing $p$ past optimal doping (the “boomerang effect”), in YBa2Cu3Ox it increases monotonically as a function of $p$ across the entire accessible range of doping [57]. Particularly notable is the fact that $T_{\theta}$ increases by $50\%$ from its value at $p=0.16$, where $T_{\rm c}=94$ K, to $p=0.19$, where $T_{\rm c}=88$ K [58] (see Figure 2. This departs markedly from the Uemura relationship, where phase stiffness is proportional to the superconducting $T_{\rm c}$ [12].
YBa2Cu3O7-δ does, in fact, exhibit an unambiguous temperature range over which phase fluctuations are visible [59, 60, 61, 62]. These experiments are consistent with 3D-XY critical fluctuations [60, 62], extend to temperatures of order 10% above $T_{\rm c}$ at optimal doping ($\delta\approx 0.07$), and for a smaller range of $T$ for overdoped YBa2Cu3O7 [59]. We conclude that phase fluctuations in overdoped YBa2Cu3O7, while much larger than in conventional superconductors (where the fluctuation regime is $\approx 10^{-9}T_{\rm c}$ [26]), are not of primary importance for determining $T_{\rm c}$.
Experimental evidence of strong phase fluctuations
[…]
Evidence that the large phase fluctuations are extrinsic
There are three classes of mechanisms that could suppress the macroscopic phase stiffness of overdoped cuprates relative to the bare Drude weight: i) strong-coupling physics, ii) pair-breaking due to scattering,[28, 29, 30] and iii) mesoscopic inhomogeneity (on scales of order the coherence length). The first mechanism is intrinsic, while the later two are extrinsic in the sense that they disappear in the limit of vanishing disorder. While all three effects may be present to some degree, we argue that extrinsic effects are dominant. Further, we suggest that mesoscopic inhomogeneity provides a more compelling explanation for several experimental phenomena, although it should be noted that the debate over the relative importance of pair breaking versus inhomogenaity is still ongoing [28, 29, 30]. We discuss this issue further in the supplementary information.
[…]
Turning to direct evidence of an inhomogeneous superconducting gap: superconducting puddles have been visualized by STM in overdoped BSCCO [75, 76, 77, 78, 79]. At $p=0.19$, superconducting gaps persist up to 120 kelvin or more in local regions with radii of a few tens of angstroms. A similar temperature scale characterizing the onset of some form of local superconducting correlations has been inferred from analysis of the Nernst effect and magnetization.[67] ARPES measurements on the same material show a particle-hole symmetric (so presumably superconducting) gap that opens at an imprecisely defined crossover temperature that remains of order 50% above $T_{c}$ for $T_{\rm c}$ as low as 70 K.[80, 81, 82, 83]
[…]
Further evidence suggesting that pairbreaking in an effective medium approximation fails to account for the suppression of $T_{\rm c}$ at the edge of the dome comes from irradiation studies on LSCO [87] and YBCO [88], and further bulk evidence of strongly electronically inhomogeneous states produced by alloy disorder in LSCO is discussed in Li et al. [89] and Tranquada et al. [32].
[…]
YBa2Cu3O7 stands alone as a material that becomes more homogeneous as it is doped from optimal to overdoped [93, 94], due to its particular mechanism of doping through copper-oxide chains. Where Tl2201, LSCO, Bi2212, and even Ca-doped YBCO generally have phase stiffnesses that decrease when moving from optimal doping to overdoped, YBa2Cu3O7’s phase stiffness increases substantially (see Figure 2). Our specific prediction is that YBa2Cu3O7 doped further using pressure or uniaxial strain would show a continued rise in superfluid phase stiffness.
Finally, we note that there are reasons to believe that the small values of $T_{\theta}$ seen in underdoped cuprates reflect distinct, more likely intrinsic physics than in the overdoped regime  [^7] This does not imply that disorder plays no role in the behavior of the underdoepd material - but it does suggest that an underlying piece of strong-correlation physics is the primary driver of the small values of $T_{\theta}$.. Specifically, there is independent evidence that not only $T_{\theta}$, but more generally the density of mobile charges drops with underdoping. For instance, the Drude weight, $D$, obtained by integrating the normal state optical conductivity up to frequencies of several hundred meV drops on the underdoped side upon approach to the Mott insuating state, while it remains large upon approaching the overdoped side of the superconducting dome; on the underdoped side the fraction of the mobile carriers that condense in the superconducting state ($\sim T_{\theta}/D$) remains nearly constant, while on the overdoped side, the drop in $T_{\theta}$ is associated with a decreasing fraction that condense.
The weak-coupling magnitude of the superconducting gap
Weak-coupling BCS theory makes a specific prediction for the ratio of the superconducting gap to the critical temperature. For a $d$-wave superconductor, this ratio depends weakly on band-structure details, of which a typical value is $\Delta/k_{\rm B}T_{\rm c}=2.14$ [95].
Experimental evidence for a weak-coupling gap
ARPES and STM measurements of the superconducting gap in BSSCO show that $\Delta/k_{\rm B}T_{\rm c}$ can be as large as 8 in the underdoped regime [96, 75]. This is indicative either of strong coupling, or of the difficult in separating the pseudogap from the superconducting gap. However, even if only the near-nodal gap[96] is assumed to be the direct consequence of pairing, this ratio grows with underdoping and reaches values in excess of 4. We will return to this later when discussing the shape of the $d$-wave gap in momentum space, but for now is enough to say that underdoped cuprates clearly do not follow the simple weak-coupling prediction for the size of the gap.
As BSSCO is hole doped past optimal doping, however, both STM and ARPES find that the gap decreases rapidly and the $\Delta_{0}/T_{c}$ ratio approaches 2 (within substantial error bars) by $p=0.24$ [83, 75, 97]. Further, while STM still shows a broad distribution of gap sizes, and correspondingly some local regions in which a gap persist to temperatures well above $T_{\rm c}$, the mean value of $\Delta_{0}/2$ and the typical temperature at which the local gap vanishes both approach $T_{\rm c}$ for $p\gtrsim 0.24$. All of these observations suggest that the strength of the superconducting gap in overdoped cuprates is roughly consistent with weak-coupling BCS theory.
Non-mean-field behavior of the gap
The persistence of superconducting gaps above $T_{\rm c}$—seen in both STM and ARPES—is inconsistent with the BCS mean-field assumption. However, as already discussed, it is natural to assign this to the intrinsic amplification of the effects of mesoscopic variations in the local environment expected in a $d$-wave superconductor with relatively short coherence length.
Absence of intertwined orders
A central feature of the cuprate phase diagram is the presence of multiple intertwined phases that all occur with comparable energy scales—antiferromagnetism, spin glass, charge density wave, and d-wave superconductivity all appear at the few-tens-of-kelvin to a couple hundred kelvin scale [98]. (If the pseudo-gap can be associated with some other form of “hidden order,”[3] then it, too, must be characterized by this same general energy scale.) The lack of separation of energy scales makes it impossible to cleanly treat any single order parameter in isolation [4, 99].
Absence of intertwined orders in overdoped cuprates
The situation is greatly simplified in the overdoped regime. There, the various non-superconducting orders clearly weaken and plausibly disappear beyond a doping that depends on the particular order under consideration, as well as on the particular cuprate. For example, charge density wave order (with correlation lengths $\xi_{cdw}\lesssim 200\AA$ ) are present in Tl2201 up to $p=0.25$, but are absent at $p=0.26$ and above. This is apparent from both resonant inlelastic X-ray scattering at $p=0.25$ and $p=0.26$ [38], as well as from the absence of Fermi surface reconstruction in quantum oscillation experiments at higher dopings [36].
Likewise in YBCO, long-range antiferromagnetism is absent above $p=0.05$, and short-range spin glass correlations are absent above $p=0.10$. Above that doping, a gap opens in the spin excitation spectrum seen in neutron scattering [100]. For fully doped YBCO ($p=0.19$), the gap in the spin excitation spectrum is of order 40 meV [101], implying that static spin structures (i.e. antiferromagnetism and related orders) and low-energy spin excitations are not relevant at these dopings. No spin order—either short or long range—has been reported in Tl2201.
Evidence of intertwined orders in overdoped cuprates
While there is no doubt the various forms of SDW and CDW stripe order seen in LSCO are strongest in underdoped LSCO, with increasingly careful studies it has become clear that weak correlations - but with much the same stripe structure - persist deep into the overdoped range of $p$, and possibly to the edge of the superconducting dome. However, careful analysis[89] strongly suggests that this order exists in the form of strongly correlated striped patches in an otherwise metallic matrix. This observation, in turn, likely reflects a similar sort of correlation enhanced amplification of the effects of otherwise innocuous mesoscopic variations in the local environment - the density-wave analogue of explanation offered for the existence of superconducting puddles.
V Concerning the “mechanism”
Once we accept the validity of the BCS perspective in the overdoped cuprates, we can work backwards from observations of the gap function to draw incomplete but clear inferences concerning the nature of the interactions responsible for the pairing. To be concrete, we focus on the ARPES results on Bi2212, where good tight-binding fits to the electronic dispersion have been obtained and for which the cleanest measurements of the superconducting gap are available.
For doping concentrations ranging from slightly underdoped to the highest level of overdoping for which data exists, and within the error bars, the $\vec{k}$ dependence of the $T\to 0$ gap function on the Fermi surface is simply given by $d(\vec{k})=(1/2)\left[\cos k_{x}-\cos k_{y}\right]$ - i.e. $f(|\vec{k}|)\approx 1$ in Eq. 6. As discussed in section II, this implies that the pairing interaction is short-ranged - primarily on nearest-neighbors.  [^8] Given the importance of this inference, it would be desirable to obtain more quantitative experimental bounds on the magnitudes of any higher gap harmonics. It would also be desirable to explore the frequency dependence of the gap function, but this is difficult without a better understanding of the full spectral function. Inverting Eq. 6, (without introducing any additional high-energy cutoff, i.e. taking $F(\epsilon)=1$) we can determine the magnitude of the pairing interaction from an empirical fit to the measured dispersion and the gap function. For the relevant range of dopings, the ARPES dispersion is well fit by
$\epsilon(\vec{k})=$ $-2t[\cos(k_{x})+\cos(k_{y})]-4t^{\prime}\cos(k_{x})\cos(k_{y})$
$-2t^{\prime\prime}[\cos(2k_{x})+\cos(2k_{y})],$ (10)
with $t=166$ meV, $t^{\prime}=-0.33t$ and $t^{\prime\prime}=0$ [102], and with a bilayer splitting of the form $2t_{\perp}d^{2}(\vec{k})$ with $t_{\perp}=0.21t$ [103]. The measured values of the gap (averaged over the bonding and anti-bonding bands) and the value we extract for the pairing strength, $\tilde{V}$, are shown in Fig. 3.
Inferences from analyzing $\Delta(\vec{k})$
  • The magnitude of the pairing interaction at optimal doping, $\tilde{V}\approx 0.6$ eV, is of the expected magnitude if one further assumes that it derives primarily from local exchange interactions. Specifically, it is well known that to next to leading order in the large $U\equiv\tilde{U}(0)$ expansion of the pure Hubbard model, one generates an effective nearest-neighbor $J\equiv\tilde{J}(1)=4t^{2}/\tilde{U}$ and $\tilde{U}(1)=-\frac{1}{4}J$. For the inferred value of $\tilde{V}=3\tilde{J}(1)-4\tilde{U}(1)=4J$, this corresponds to $J\approx 0.15$ eV, in reasonable agreement[104] with values inferred in undoped antiferromagnet cuprates.[105] A more ambitious analysis of the pairing mechanism in optimally doped YBCO has been carried through in Dahm et al. [106] and Tacon et al. [107] using the measured magnetic structure factor as the bosonic “glue” in the Migdal-Eliashberg version of the gap equation. This approach gives a mean-field $T_{c}$ roughly twice the actual $T_{c}$, consistent with the conjecture that both pairing and phase fluctuations are of comparable importance in this crossover regime. Significantly, a relatively wide range of energies (of order the width of the paramagnon dispersion, $\Delta E\sim 2J\approx 0.3$eV) contribute to the pairing interaction. This implies that the frequency dependence of the interactions can be neglected at lower energies (including energies comparable to the superconducting gap), i.e. the results should be equivalent to those obtained with a BCS approach.
  • To the extent that $J$ is an extension of the local physics of the undoped antiferromagnet, one would expect it to decrease with large enough doping. The smooth but none-the-less pronounced decrease of $\tilde{V}$ with increasing $p$ thus provides additional suggestive evidence that it is largely magnetic in origin. Comparing this expectation with direct probes of the doping dependence of the magnetic fluctuations is somewhat complicated. X-ray measurements[108, 109, 110] suggest that there is relatively little decrease in the strength of the magnetic fluctuations over much of the Brillouin zone (especially near $\vec{Q}=(\pi,0)$). However, neutron scattering[111, 112, 89] studies have found that in the important[113] range of $\vec{Q}$ near $(\pi,\pi)$, there is indeed a pronounced drop in magnetic intensity with overdoping.
  • The fact that the measured $\vec{k}$-dependence of the gap is well approximated by $d(\vec{k})$ implies that the effective interaction remains short-range as low as optimal doping. This precludes any theory that identifies optimal doping with the point at which the pairing arises primarily from the exchange of any familiar quantum-critical collective mode. (Note, this does not imply that there are no quantum critical points under the superconducting dome - given the plethora of intertwined orders, such quantum critical points may well arise in some circumstances. [^9] For instance, the observation[114] of an apparently divergent electron effective mass seen in quantum oscillations in YBCO as $p$ approach a critical value $\sim p_{opt}$ is strong evidence of the existence of a quantum critical point in the “normal” state obtained when superconductivity is quenched with a high magnetic field. However, to the extent that $\Delta(\vec{k})\propto d(\vec{k})$, we conclude that no quantum critical mode is the prime driver of superconductivity.)
    From this perspective, it is interesting to compare the $\vec{k}$-dependence of the gap in the hole-doped cuprates with the corresponding (not nearly as well documented) behavior of the electron-doped cuprates. In the electron doped materials, even at optimal doping there is direct evidence of antiferromagnetic fluctuations with a correlation length large compared to one lattice constant. To the extent that these fluctuations play a significant role in the pairing interaction, one would expect this to lead to a much more structured $\Delta(\vec{k})$. Indeed, suggestive evidence of a non-monotonic dependence of the gap along the Fermi surface has been observed - see Ref. [35] for a review.
Additional Considerations
  • It is important to acknowledge that, at short-distances and high energy scales, it follows from the low electron density (large value of $r_{s}$) that the cuprates are more strongly interacting than conventional metals. The above treatment of the mechanism of superconductivity in overdoped cuprates does not contradict this observation. The essential assumption we are making is that when high energy degrees of freedom are integrated out, the quasi-particles in the resulting low-energy effective Hamiltonian are sufficiently weakly interacting for a Fermi-liquid description to be possible; both the parameters that characterize the quasi-particle dispersion relation in Eq. 10 and the effective pairing interaction, $\tilde{V}$, should thus be interpreted as effective parameters.
  • One can legitimately question the self-consistency of this analysis, given that we have obtained values of the pairing interaction, $\tilde{V}$, that are not small, i.e $\tilde{V}\rho(E_{F})\sim 1$ at optimal doping, dropping to $\tilde{V}\rho(E_{F})\sim 0.5$ at the highest doping we analyze in Figure 3. However, the same can be said of the Migdal-Eliashberg treatment of conventional SCs such as Pb, where in fact the dimensionless strength of the electron-pohonon coupling, $\lambda\sim 1$[115].
Figure 3: Effective pairing interaction inferred from ARPES data for Bi2212 (from [83]): The yellow points are the measured values of $\Delta_{0}$ and the blue dots are the expected BCS magnitude of $\Delta_{0}=2.14k_{B}T_{c}$ for the measured $T_{c}$. The red points are the values of the pairing interaction, $\tilde{V}$, obtained by inverting Eq. 7 using the measured values of $\Delta_{0}$ and of the normal-state dispersion. The grey and red lines are simple fits to the existing data that allow it to be smoothly extrapolated to higher values of doping, for which no data exists. The purple dots are the values of $\Delta_{0}$ that would result from values of $\tilde{V}$ on the red line. The inset is a blow-up of the range of doping near the critical point at which the extrapolated $\tilde{V}$ would change sign.
VI Falsifiable Consequences
While it is not possible to have a disorder-free material, it is possible to identify circumstances in which the amount of disorder is significantly suppressed. The principal empirical test of the perspective presented here is that deviations from the predictions of Fermi liquid theory for the low energy properties of the normal state, and of BCS mean-field theory for the superconducting state, should tend to be decreasingly pronounced in overdoped cuprates the weaker the disorder. Below, we list some of the experimental properties that could serve as particularly good indicators of such trends, and where possible, we mention some existing data that constitutes encouraging evidence of the expected tendencies.
  • The linear drop of $T_{c}\sim(p_{max}-p)$ to zero at a critical doping $p_{max}$ should get increasingly rounded (convex) and $p_{max}$ is likely to increase (see Figure 3). The fact that $T_{\rm c}$ persists up to at least $p=0.31$ in Tl2201 [36], compared to $p=0.27$ in the comparatively more disordered LSCO, is suggestive of this, but a finer doping dependence is needed
  • The superfluid density should grow to be an increasingly large fraction of the total Drude weight. Specifically, one should observe an increasing tendency toward the edge of the superconducting dome to realize the familiar situation in which $T_{\theta}\gg T_{c}$. The data in slightly-overdoped YBa2Cu3Ox is suggestive that this may be the case in the cleanest cuprates, but higher dopings are needed.
  • The zero temperature gap function should approach the BCS value in magnitude, $\Delta_{0}\approx 2T_{c}$. Moreover, spectroscopic evidence for a gap that persists for $T>T_{c}$, or beyond the end of the superconducting dome, $p>p_{max}$, should occur over decreasing ranges of $T$ and $p$. The existing data in Bi2212 is suggestive that this is the case (see Figure 3), but higher dopings are needed.
  • The specific heat should show increasingly clear mean-field features with decreasing disorder, including a mean-field jump in the specific heat at $T_{c}$ and a strong supression of the residual $C/T$ in the $T\to 0$ limit. There should be a similar drop in the residual Drude response seen in terahertz optical conducitivity for $\omega\ll\Delta_{0}$ and $T\ll T_{c}$.
  • The $T$-linear term in the resistivity seen in overdoped materials when superconductivity is suppressed with a magnetic field should become increasingly small with decreasing disorder.
YBa2Cu3O7 is somewhat overdoped and is stoichiometric, meaning that intrinsic alloy disorder is absent; indeed, various experimental probes appear to corroborate the expectation that available samples of YBa2Cu3O7 are tangibly less electronically inhomogeneous than other overdoped materials. For the future, it seems that the doped hole concentration in the Cu-O planes of this material can be increased with application of pressure, so this may be the ideal material to test the present ideas.  [^10] It is also observed that in multi-layer cuprates, the outer-planes screen the disorder intrinsic to the charge-reservoir layers, meaning that disorder effects are reduced on the inner planes. Unfortunately, for present purposes, the inner layers tend to be underdoped - often so much so that they exhibit long-range antiferromagnetic order coexisting with the superconductivity (i.e. it is a distinct phase of matter).
Acknowledgments
We would like to acknowledge extremely useful discussions with P. Armitage, B. Atkinson, M. Beasley, I. Bozovic, S. Brown, S. Chakravarty, A. Chubukov, R. Greene, H. Guo, S. Hartnoll, P. Johnson, M.-H. Julien, B. Keimer, R.B. Laughlin, A. Millis, M. Norman, S. das Sarma, J. Schmalian, M. Sulangi, S. Uchida, T. Valla, and H-H. Wen and particularly detailed and insightful feedback from Y. He, J. Tranquada, P. Hirschfeld, and D. Broun. SAK was supported in part by NSF-BSF award DMR-2310312. BJR was supported in part by NSF award DMR-2428169 and the Canadian Institute for Advanced Research.
Author Contributions
B.J.R and S.A.K contributed equally to all aspects of the project.
Competing interests
The authors declare no competing interests.
Data and Code availability
All data are available from the authors upon request.